A car running at a speed of 72 km/h is to be stopped in exactly 10 seconds. Find the uniform deceleration needed.

1
Expert's answer
2021-12-15T11:19:38-0500

Explanations & Calculations


  • The question mentions a uniform deceleration. For uniform acceleration/ deceleration 4 equations of motions can be used appropriately.

"\qquad\qquad
\begin{aligned}
\small v=u+at,\, v^2=u^2+2as, \,s=ut+\frac{1}{2}at^2,\, s=\frac{(v+u)t}{2}
\end{aligned}"

  • The appropriate equation to be used for this question is the simplest one which is "\small v=u+at" as the initial speed ("\small 72kmh^{-1}" ), final speed and the time are given.
  • To begin with the calculations, you need to convert the given initial speed into "\small ms^{-1}" as the time is given in seconds.

"\qquad\qquad
\begin{aligned}
\small 72\,kmh^{-1}&=\small \frac{72\times1000\,m}{3600\,s}\\
&=\small 20\,ms^{-1}
\end{aligned}"

  • Now you can try substituting the values to the equation and getting the answer.
  • The answer should come with a negative sign as the finding is the deceleration, not the acceleration.